# C2P: Featuring Large Language Models with Causal Reasoning

### v2

@@ Abstract
Causal reasoning is the primary bottleneck that Large Language Models (LLMs) must overcome to attain human-level intelligence. To address this, we introduce the Causal Chain of Prompting (C2P) as the first reasoning framework that equips current LLMs with causal reasoning capabilities. C2P operates autonomously, avoiding reliance on external tools or modules during both the causal learning and reasoning phases, and can be seamlessly implemented during the training or fine-tuning of LLMs. Experimental results across various benchmark datasets demonstrate a significant improvement in causal learning and subsequent reasoning accuracy of LLMs. We illustrate how C2P enhances LLMs’ ability to causally reason in real-world scenarios, addressing complex problems in fields such as healthcare, medicine, economics, education, social sciences, environmental science, and marketing. With few-shot learning, GPT-4 Turbo using C2P with as few as six examples achieves significant performance improvements, boasting over a $33\%$ increase in reasoning accuracy over the most state-of-the-art LLMs, which perform nearly randomly in similar circumstances. This demonstrates the transformative potential of integrating C2P into LLM training or fine-tuning processes, thereby empowering these models with advanced causal reasoning capabilities.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have impacted existing AI paradigms and heightened expectations regarding AI’s capabilities (Achiam et al., 2023; Brown et al., 2020).
Figure 1: Example of the standard prompting vs few-shot learned GPT-4 with C2P in open problem in astrophysics (Pasquato et al., 2023)
Despite significant architectural differences in LLMs, they generally produce outputs based on the most likely results learned from vast amounts of training data (Vaswani et al., 2017). This enables them to acquire extensive knowledge ranging from common sense to specialized domains such as mathematics and science (Jiralerspong et al., 2024). Despite this, the inefficiency of LLMs in addressing causal reasoning questions remains their primary bottleneck and simple tasks can completely break down reasoning in state-of-the-art LLMs (Nezhurina et al., 2024). Additionally, studies such as Kalai & Vempala (2023); Xu et al. (2024) have demonstrated that, despite the training data containing numerous examples of interventions, outcomes, and explanations, as well as similar tasks, hallucinatory responses persist and there is a lack of causal reasoning capability. As a result, while they may talk causality, they are not causal (Zečević et al., 2023). This deficiency represents a fundamental drawback of LLMs as AI systems compared to human intelligence, which goes beyond mere correlations and depends on causal relationships for decision-making (Penn & Povinelli, 2007; Anwar et al., 2024).
@@
Recently, answering cause and effect questions with LLMs has gained extensive interest (Shin et al., 2020; Jin et al., 2023b; Ashwani et al., 2024). To address this issue, it is important to note that LLMs have been utilized in conjunction with external tools to extract causal structures, as demonstrated in (Jiralerspong et al., 2024). However, their architectures lack specialized modules specifically designed to enhance the understanding of cause-and-effect relationships within their outputs (Wang et al., 2023; Imani et al., 2023). Aside from studies that reason causally based on the knowledge that is already in their training data (Petroni et al., 2019; Jiang et al., 2020) and the ones that use LLMs in causality (Kıcıman et al., 2023; Zhang et al., 2023; Feder et al., 2024; Khatibi et al., 2024), chain-of-thought prompting is presented in (Wei et al., 2022) as one of the initial attempts in enhancing the reasoning in LLMs that shows improvement based on the data of the given query. However, LLMs still struggle with rigorous numerical and abstract reasoning among many other tasks (Xu et al., 2023). For example, a recent work, Causal Reasoning Assessment Benchmark (CRAB, Romanou et al., 2023), is designed to evaluate the causal understanding of events in real-world narratives. This study demonstrated that most systems perform poorly in identifying cause-and-effect tasks. Similarly, in (Jin et al., 2023b), the CORR2CAUSE dataset is introduced and demonstrated that current models often perform no better than random chance when tasked with causal questions. Following that, in (Jin et al., 2023a), the CLADDER dataset is introduced to asses Average Treatment Effects with LLMs and it is demonstrated that these models struggle with causal tasks. In their study, with the implementation of the proposed framework, CAUSALCoT, progress has been made in evaluating the average treatment effect where LLMs are provided with a collection of causal graphs and various types of queries (associational, interventional, and counterfactual), such as those included in the CLADDER dataset. More recently, in (Ashwani et al., 2024), a novel architecture called the Context-Aware Reasoning Enhancement with Counterfactual Analysis (CARE-CA) framework is presented to enhance causal reasoning and explainability. Their proposed framework incorporates an external explicit causal detection module with ConceptNet (Speer et al., 2017) and counterfactual statements, as well as implicit causal detection through LLMs, showing progress in causal reasoning in short and simple queries. Several other works at the intersection of causal inference and LLMs are discussed in an extensive survey by Liu et al. (2024).
@@
As argued by Pearl (1995), causal Directed Acyclic Graphs (DAGs), along with d-separation, allow for the investigation of cause-and-effect relationships without relying on structural equation models in computational studies. Inspired by Pearl’s foundational work, we propose a novel framework, named Causal Chain of Prompting (C2P), to address the inefficiencies of LLMs in handling causal queries (see Figure 1 as an example). We show that by identifying the adjacency matrix, equivalent to and instead of the causal DAG in Pearl’s work, the causal relationships among the variables in the premise’s cause-and-effect relationships can be effectively reasoned within the context of language models. Contrary to the existing research in addressing weaknesses of LLMs in causal reasoning, the C2P operates autonomously, avoiding reliance on external tools or modules during both the learning and reasoning phases to answer causal questions. Additionally, C2P can easily be implemented in the training or fine-tuning process of LLMs. C2P includes five simple main sub-tasks, as follows: (1) Prompting for extracting the random variables from the provided data. (2) Prompting for extracting all conditional and unconditional relations and cause and effect relations specifically mentioned among the random variables. (3) Prompting to create the initial adjacency matrix with values 1 for all elements except the diagonal elements and the ones corresponding to effect-cause relations (the cause and effects elements are also one). (4) Prompting of the conditional and unconditional independencies and identification of the colliders, step by step, to extract the causal adjacency matrix; and, (5) Prompting for reasoning questions or hypotheses. To evaluate the accuracy and reliability of implementing the C2P on LLMs, we initially assess it using publicly available synthetic datasets such as (Jin et al., 2023b). Subsequently, we evaluate it in more realistic and complex scenarios in real-world problems presented in (Pearl & Mackenzie, 2018) and (Pasquato et al., 2023). Moreover, we present results on few-shot learning with C2P on both synthetics and realistic scenarios.
@@
In this work, we present several important contributions to facilitate causal reasoning in language models. Concretely,
@@
    We introduce the C2P framework as the first reasoning framework to equip LLMs with the ability of causal reasoning within real-world scenarios, without relying on external tools.
@@
    Through extensive experiments with our framework, we demonstrate a significant improvement of LLMs in causal reasoning in various benchmarks. Additionally, we examine the performance of the C2P framework on more complex and real-world scenarios in various domains.
@@
    We perform few-shot learning experiments in section 4 with GPT-4 Turbo using our C2P framework and show how integrating C2P during the training or fine-tuning of LLMs can revolutionize existing models. This approach equips them with causal reasoning capabilities akin to the transformative impact of ‘Chain of Tought’ (Wei et al., 2022) on LLMs, as discussed by Chung et al. (2024).
The codes for sections 3 and 4 are publicly available at https://anonymous.4open.science/r/C2P-5C2A/.
@@ 2 Preliminaries on Causal Learning and Reasoning
To lay the groundwork for our framework, understanding the fundamental concept of cause and effect is paramount. A foundational and widely acknowledged principle in causality is the one proposed by Reichenbach, which posits the following:
Principle 1 (Common cause principle (Reichenbach, 1991))
[…]
Markov Property: The Markov property in a DAG $G$ states that each node $X_{i}$ is conditionally independent of its non-descendants given its parents, namely $X_{i}\perp\!\!\!\perp\text{NonDe}(X_{i})|\text{Pa}(X_{i})$, where $\text{NonDe}(X_{i})$ denotes the non-descendants of $X_{i}$ excluding itself, and $\text{Pa}(X_{i})$ denotes the parents of $X_{i}$. Using the Markov property, we can factorize the joint distribution of all the nodes in the graph into $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i}|\text{Pa}(X_{i}))$. To infer the causal graph from probability distributions, a common assumption is faithfulness, namely the validity to infer all the d-separation sets in the graph from the independence relations in the probability distribution. In our work, we also take this broadly taken assumption which holds for most real-world scenarios.
@@
Directed Acyclic graph: A graph $G$ is called a partially directed acyclic graph (PDAG) if there is no directed cycle, that is, if there is no pair $(j,k)$ with directed paths from $j$ to $k$ and from k to $j$. $G$ is called a directed acyclic graph (DAG) if it is a PDAG and all edges are directed.
@@
Markov Equivalence of Graphs: We denote two DAGs as Markov equivalent if they induce the same joint distribution $P(\mathbf{X})$. The set of DAGs that are Markov equivalent to each other is called a Markov equivalence class (MEC). Causal graphs in the same MEC can be easily identified since they have the same skeleton (i.e., undirected edges) and V-structures (i.e., structures in the form of $X_{1}\rightarrow X_{2}\leftarrow X_{3}$ where $X_{1}$ and $X_{3}$ are not connected).
@@
The be able to perform any level of causal reasoning, the Ladder of Causation is proposed with three main rungs, “seeing”, “doing”, and “imagining”, in (Pearl & Mackenzie, 2018). To fulfill all rungs of the ladder, more information on the causal structure is needed, which can be mostly provided with structural causal models (SCM) (Assuming all the assumptions are satisfied) (Bareinboim et al., 2022). The current causal discovery methods are divided into two groups, constraint-based algorithm such as the PC algorithm, which has a PDAG as an output, that represents the MEC of the true underlying graph and is the best result such methods can extract. Second is score-based methods such as NOTEARS (Zheng et al., 2018), GOLEM (Ng et al., 2020), and DAGMA (Bello et al., 2022), among many others, that extract a DAG that mostly fits the data and demonstrated very high accuracy in extracting Bayesian networks. Score-based methods require solving an optimization problem based on the numerical values of variables, rendering them impractical for implementation in LLM models.
@@
PC algorithm: Following these materials, the PC algorithm is developed based on Reichenbach’s common cause principle and Markov property as follows for a causal mechanism with $p$ random variables (Spirtes et al., 2001);
@@
    For each pair of variables $X_{1},X_{2}$ having an edge between them, and for each pair of variables $\{X_{3},X_{4}\}$ with edges both connected to $X_{1}$ or both connected to $B$, eliminate the edge between $X_{1}$ and $X_{2}$ if $X_{1}\perp\!\!\!\perp X_{2}|\{X_{3},X_{4}\}$.
@@
    For each triple of variables $(X_{1},X_{2},X_{3})$ such that $X_{1}$ and $X_{2}$ are adjacent, $X_{2}$ and $X_{3}$ are adjacent, and $A$ and $C$ are not adjacent, orient the edges $X_{1}-X_{2}-X_{3}$ as $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, if $X_{2}$ was not in the set conditioning on which $X_{1}$ and $X_{3}$ became independent and the edge between them was accordingly eliminated. We call such a triple of variables a V-structure.
@@
d-separation and do calculus (Pearl, 1995): do-calculus, developed by Judea Pearl and others, is a set of rules used to transform and manipulate causal expressions within causal diagrams (or graphical models). It is a generalization of do-calculus, which is itself a formal tool used to reason in causal relationships from a mixture of experimental and observational data. do-calculus consists of three main rules that enable one to rewrite expressions involving interventions (typically represented as $do(x)$, indicating an intervention to set variable $X$ to value $x$):
  1. 1.
  2. 2.
  3. 3.
These rules are crucial for determining the identifiability of causal effects from data, allowing researchers to reason in causal relationships from a combination of experimental and observational data.
d-Separation is a criterion used in the analysis of Bayesian networks to determine whether a set of variables $X_{1}$ is independent of another set of variables $X_{2}$, given a third set of variables $X_{3}$. This concept is foundational in understanding the flow of causal influence in graphical models and helps in deciding whether a path between two variables is "blocked" or not by conditioning on other variables. The basic rule of d-separation is: A path between two variables is blocked if it includes an intermediate variable that is a collider and is not conditioned on or a non-collider that is conditioned on. Here, a collider is a variable that has inward arrows from two other nodes (i.e., $X_{1}\rightarrow X_{2}\leftarrow X_{3}$), whereas a non-collider does not meet this criterion.
3 Developing Causal Chain of Prompting

### v4

@@ Abstract
Causal reasoning is one of the primary bottlenecks that Large Language Models (LLMs) must overcome to attain human-level intelligence. Recent studies indicate that LLMs display near-random performance on reasoning tasks. To address this, we introduce the Causal Chain of Prompting ($\text{C}^{2}\text{P}$), a reasoning framework that aims to equip current LLMs with causal reasoning capabilities as the first framework of its kind operating autonomously without relying on external tools or modules during both the causal learning and reasoning phases. To evaluate the performance of $\text{C}^{2}\text{P}$, we first demonstrate that reasoning accuracy improved by over $30.7\%$ and $25.9\%$ for GPT-4 Turbo and LLaMA 3.1, respectively, when using our framework, compared to the same models without $\text{C}^{2}\text{P}$ on a synthetic benchmark dataset. Then, using few-shot learning of the same LLMs with $\text{C}^{2}\text{P}$, the reasoning accuracy increased by more than $20.05\%$ and $20.89\%$, respectively, with as few as ten examples, compared to the corresponding LLMs without $\text{C}^{2}\text{P}$ on the same dataset. To evaluate $\text{C}^{2}\text{P}$ in realistic scenarios, we utilized another benchmark dataset containing natural stories across various fields, including healthcare, medicine, economics, education, social sciences, environmental science, and marketing. The results show improved reasoning when $\text{C}^{2}\text{P}$ is applied, compared to cases where our framework is not used, which often leads to random and hallucinated responses. By showing the improved performance of few-shot learned GPT-4 Turbo and LLaMA 3.1 with $\text{C}^{2}\text{P}$, we demonstrate the generalizability of our framework.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have impacted existing AI paradigms and raised expectations regarding AI’s capabilities (Achiam et al., 2023; Brown et al., 2020).
Figure 1: The 5 successive subtasks of $\text{C}^{2}\text{P}$ applied to Black Holes and their Host Galexis example presented in (Pasquato et al., 2023). Central Density(CD), Black Hole Mass(BHM), Velocity Density(VD), Bulge Stellar Mass(BSM), Effective Radius(ER)
LLMs generally produce outputs based on the most likely results learned from vast amounts of training data (Vaswani et al., 2017). This enables them to acquire extensive knowledge, ranging from common sense to specialized domains such as mathematics and science (Jiralerspong et al., 2024). Numerous examples of interventions, outcomes, and explanations are included in their training of LLMs to reduce hallucinations and improve reasoning. However, despite significant architectural differences, hallucinatory responses still occur and true causal reasoning remains lacking (Kalai & Vempala, 2023; Xu et al., 2024). As a result, although models may appear to reason causally, they do not engage in a genuine causal reasoning process (Zečević et al., 2023). This deficiency represents a fundamental limitation of LLMs as AI systems compared to human intelligence, which is based on causal reasoning rather than simple associations for decision making (Penn & Povinelli, 2007; Anwar et al., 2024). Judea Pearl introduced "The Ladder of Causation" in (Pearl & Mackenzie, 2018) that easily addresses the reason for this deficiency. This ladder includes three main levels. At the first level, association, only patterns and dependencies are identified, and why things are related cannot be answered. The second level, intervention, involves understanding cause-and-effect by predicting outcomes of actions or changes, asking "What happens if I do this?" The highest level, counterfactuals, deals with imagining alternative scenarios, asking "What would have happened if…?". This framework illustrates how reasoning evolves from pattern recognition to understanding and predicting causal relationships, highlighting the differences at each level.
@@
Recently, studies on the causal reasoning capabilities of LLMs have garnered significant interest. Most of these focus on evaluating the reasoning abilities of LLMs, while a smaller number aim to enhance this feature. Additionally, some studies explore the application of LLMs in the field of causality.(Kıcıman et al., 2023; Zhang et al., 2023; Feder et al., 2024; Khatibi et al., 2024). The inefficiency of LLMs in reasoning has been extensively studied from various perspectives, as demonstrated in works such as (Xu et al., 2023; Romanou et al., 2023; Jin et al., 2023a; Liu et al., 2024a; Hobbhahn et al., 2022). Similarly, Jin et al. (2023b) introduces the CORR2CAUSE dataset, revealing that current models often perform no better than random chance when answering causal questions. Most recently, simple tasks have been shown to completely break down the reasoning abilities of state-of-the-art LLMs (Nezhurina et al., 2024). Additionally, studies such as (Petroni et al., 2019; Jiang et al., 2020) aimed to reason causally based on the knowledge already present in the training data of LLMs, which is why LLMs are referred to as "causal parrots" in (Zečević et al., 2023). The reason for these behaviors is discussed in (Wang et al., 2023; Imani et al., 2023). As one of the initial attempts to enhance reasoning in LLMs, chain-of-thought prompting is presented in (Wei et al., 2022b), showing improvement based on the data from the given query. Despite CoT’s potential to enhance reasoning performance, the analysis in Bao et al. (2024) reveals a pattern: correct answers often follow incorrect CoTs, and vice versa. As another approach, LLMs have been utilized in conjunction with external tools to extract causal structures, as demonstrated in (Jiralerspong et al., 2024). More recently, in (Ashwani et al., 2024), a novel architecture called the Context-Aware Reasoning Enhancement with Counterfactual Analysis (CARE-CA) framework is presented to enhance causal reasoning and explainability. Their proposed framework incorporates an external explicit causal detection module with ConceptNet (Speer et al., 2017) and counterfactual statements, as well as implicit causal detection through LLMs, showing progress in causal reasoning in short and simple queries. Several other works at the intersection of causal inference and LLMs are discussed in an extensive survey by Liu et al. (2024b). The main drawbacks of the existing frameworks aiming to equip LLMs with reasoning are their reliance on external modules, the need for extensive information to function, and their very low accuracy.
@@
In this paper, we propose the Causal Chain of Prompting ($\text{C}^{2}\text{P}$), a reasoning framework for LLMs designed to enhance reasoning skills by climbing the causality ladder and answering reasoning questions. Unlike existing methods, $\text{C}^{2}\text{P}$ operates autonomously, without relying on external tools or modules during the learning and reasoning phases. It can be easily implemented in the few-shot to improve reasoning in causal questions. Importantly, $\text{C}^{2}\text{P}$ extracts the causal relation based on associations mentioned in a given premise. $\text{C}^{2}\text{P}$ is inspired by Pearl’s foundational work, which argues that causal Directed Acyclic Graphs (DAGs), along with d-separation, enable the investigation of cause-and-effect relationships without relying on structural equation models in computational studies (Pearl, 1995). Based on this, we demonstrate that by identifying the adjacency matrix of causal relationships among variables in the premise—similar to, but distinct from, the causal DAG in Pearl’s framework—a reasoning query can be effectively answered. This framework includes five simple main subtasks, as follows: (1) Prompting to extract random variables from the provided data. (2) Prompting to extract all conditional and unconditional associations, as well as cause-and-effect relations specifically mentioned among the random variables. (3) Prompting to create the initial adjacency matrix with values of 1 for all elements except the diagonal elements and those corresponding to effect-cause relations (the cause-and-effect elements are also set to 1). (4) Prompting conditional and unconditional independencies and identification of colliders, step by step, to extract the causal adjacency matrix. (5) Prompting for reasoning questions or hypotheses. (see Fig. 1). To evaluate the accuracy and reliability of implementing the $\text{C}^{2}\text{P}$ on LLMs, we initially assess it using publicly available benchmark synthetic and "Natural Story" datasets, such as those in (Jin et al., 2023b). To demonstrate the practical applicability of $\text{C}^{2}\text{P}$, we present the results of few-shot learned LLMs with $\text{C}^{2}\text{P}$ in both synthetic and real-world scenarios. Subsequently, we evaluate it in more realistic and complex scenarios found in real-world problems presented in (Pasquato et al., 2023).
@@
In this work, we present several contributions to facilitate causal reasoning in large language models. Concretely,
@@
    We introduce the $\text{C}^{2}\text{P}$ framework as the first reasoning framework to equip LLMs with causal reasoning within real-world scenarios, without relying on external tools.
@@
    Through extensive experiments with our framework, we demonstrate significant improvements in the causal reasoning abilities of LLMs across various benchmark datasets, as well as in more complex, real-world scenarios across multiple domains.
@@
    We demonstrate the generality of $\text{C}^{2}\text{P}$ by implementing few-shot learning for both GPT-4 Turbo and LlaMA 3.1, obtaining significant improvements in reasoning.
The code is publicly available at https://github.com/abmbagheri/c2p-Featuring-Large-Language-Models-with-Causal-Reasoning.git.
@@ 2 Preliminaries on Causal Learning and Reasoning
To lay the groundwork for our framework, understanding the fundamental concept of cause and effect is paramount. A widely acknowledged principle in causality is the principle proposed by Reichenbach, which posits the following.
Common cause principle (Reichenbach, 1991) If two random variables $X_{1}$ and $X_{2}$ are statistically dependent, i.e., $X_{1}\not\!\perp\!\!\!\perp X_{2}$, then there exists a third variable $X_{3}$ that causally influences both. (As a special case, $X_{3}$ may coincide with either $X_{1}$ or $X_{2}$.) Furthermore, this variable $X_{3}$ screens $X_{1}$ and $X_{2}$ from each other in the sense that given $X_{3}$, they become independent, $X_{1}\perp\!\!\!\perp X_{2}|X_{3}$.
do calculus (Pearl, 1995): do-calculus, developed by Judea Pearl et al., is a set of rules used to transform and manipulate causal expressions within causal diagrams (or graphical models). do-calculus is a formal tool used to reason in causal relationships from a mixture of experimental and observational data. do-calculus consists of three main rules that allow one to rewrite expressions involving interventions (typically represented as $do(x)$, indicating an intervention to set variable $X$ to value $x$). These rules are crucial for determining the identifiability of causal effects from data, allowing researchers to reason about causal relationships using a combination of experimental and observational data. The rules are provided in Appendix A.1.
d-separation (Pearl, 1995). d-Separation is a criterion used in Bayesian network analysis to determine whether a set of variables $X_{1}$ is independent of another set of variables $X_{2}$, given a third set of variables $X_{3}$. This concept is foundational in understanding the flow of causal effects in graphical models and helps in deciding whether a path between two variables is "blocked" or not by conditioning on other variables. According to the d-separation, a path between two variables is blocked if it includes an intermediate variable that is a collider and is not conditioned on or a non-collider that is conditioned on. Here, a collider is a variable that has arrows inward from two other nodes (i.e., $X_{1}\rightarrow X_{2}\leftarrow X_{3}$), whereas a non-collider does not meet this criterion.
@@
Directed Acyclic Graphs. A graph $G$ is called a Partially Directed Acyclic Graph (PDAG) if there is no directed cycle, that is, if there is no pair $(X_{j},X_{k})$ with directed paths from $X_{j}$ to $X_{k}$ and from k to $j$. $G$ is called a Directed Acyclic Graph (DAG) if it is a PDAG and all edges are directed.
Markov Property. The Markov property in a DAG $G$ states that each node $X_{i}$ is conditionally independent of its non-descendants, given its parents. In other words, $X_{i}\perp\!\!\!\perp\text{NonDe}(X_{i})|\text{Pa}(X_{i})$, where $\text{NonDe}(X_{i})$ represents the non-descendants of $X_{i}$, excluding itself, and $\text{Pa}(X_{i})$ represents the parents of $X_{i}$. It helps factorize the distribution of all graph nodes as $P(X_{1},\ldots,X_{N})=\prod_{i=1}^{N}P(X_{i}|\text{Pa}(X_{i}))$.
Faithfulness. This assumption ensures that all the d-separation sets in the graph can be inferred from the independence relations in the distribution. In the sequel, we assume faithfulness, a widely used assumption in causal discovery (Spirtes et al., 2001).
@@
Markov Equivalence of Graphs. Two DAGs are Markov equivalent if they generate the same joint distribution, $P(\mathbf{X})$. A set of DAGs that are Markov equivalent is referred to as a Markov equivalence class (MEC). Causal graphs within the same MEC are easily recognizable because they share the same skeleton (i.e., the same undirected edges) and the same V-structures (i.e., configurations like $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, where $X_{1}$ and $X_{3}$ are not directly connected).
Ladder of Causation (Pearl & Mackenzie, 2018). To perform any level of causal reasoning, the Ladder of Causation proposes three main levels: “seeing or association,” “doing or intervention,” and “imagining or counterfactual”. The questions that can be answered in Association (Seeing) are mostly similar to "What is happening?" This is the most basic level where we observe patterns and correlations between variables. At this level, we can only say that two things are related or tend to occur together, but we cannot explain why. Machine learning models and statistical methods that rely on pattern recognition (like LLMs) often operate at this level. In the second level, Intervention (Doing), questions such as "What happens if I do something?" can be addressed. At this level, we can go beyond mere association and ask about the effect of an action or intervention. This requires understanding the cause-and-effect relationships. To reach the third level of the ladder, more information is needed on the causal structure, which can mainly be provided with structural causal models (SCMs), assuming that all assumptions are satisfied (Bareinboim et al., 2022). However, studies such as (Spirtes et al., 2001) have shown that Level 2 can be reached (up to equivalence classes) using only observational data.
@@
Note that the current causal discovery methods are primarily divided into two groups: (i) Constraint-based algorithms such as the PC algorithm (Spirtes et al., 2001), which has a PDAG as an output, which represents the MEC of the true underlying graph and is the best outcome that these methods can achieve; (ii) Score-based methods such as GES (Chickering, 2002), NOTEARS (Zheng et al., 2018), GOLEM (Ng et al., 2020), DAGMA (Bello et al., 2022), and TOPO (Deng et al., 2023), among many others, which extract a DAG that mostly fits the data and demonstrated high accuracy in extracting Bayesian networks. Nevertheless, score-based methods require the solution of a numerical optimization problem, making them difficult to integrate with LLMs.
@@
The PC algorithm (Spirtes et al., 2001). The PC algorithm is developed based on Reichenbach’s common cause principle and the Markov property, its steps can be described as follows:
@@
    For each pair of variables $X_{1},X_{2}$ having an edge between them, and for each pair of variables $\{X_{3},X_{4}\}$ with edges both connected to $X_{1}$ or both connected to $X_{2}$, eliminate the edge between $X_{1}$ and $X_{2}$ if $X_{1}\perp\!\!\!\perp X_{2}|\{X_{3},X_{4}\}$.
@@
    For each triple of variables $(X_{1},X_{2},X_{3})$ such that $X_{1}$ and $X_{2}$ are adjacent, $X_{2}$ and $X_{3}$ are adjacent, and $X_{1}$ and $X_{3}$ are not adjacent, orient the edges $X_{1}-X_{2}-X_{3}$ as $X_{1}\rightarrow X_{2}\leftarrow X_{3}$, if $X_{2}$ was not in the set conditioning on which $X_{1}$ and $X_{3}$ became independent and the edge between them was accordingly eliminated. We call such a triple of variables a V-structure.
@@
3 The Causal Chain of Prompting